### v1

# Aleph Filter: To Infinity in Constant Time

Abstract.
Filter data structures are widely used in various areas of computer science to answer approximate set-membership queries. In many applications, the data grows dynamically, requiring their filters to expand along with the data that they represent. However, existing methods for expanding filters cannot maintain stable performance, memory footprint, and false positive rate at the same time. We address this problem with Aleph Filter, which makes the following contributions. (1) It supports all operations (insertions, queries, deletes, etc.) in constant time, no matter how much the data grows. (2) Given any rough estimate of how much the data will ultimately grow, Aleph Filter provides far superior memory vs. false positive rate trade-offs, even if the estimate is off by orders of magnitude.
1. Introduction
Filters. A filter is a compact data structure that represents keys in a set and answers approximate set-membership queries. It cannot return false negatives, but it returns a false positive with a probability that depends on the number of bits used on average to represent each key. A filter is typically stored at a higher layer of the memory hierarchy (e.g., DRAM) than the base data, which typically resides in storage (i.e., disk or SSD) and/or over a network. If a filter returns a negative, the sought-after key is guaranteed not to exist, meaning that the full data set does not need to be searched. In this way, filters eliminate redundant storage accesses and network hops to improve a system’s performance.
From Bloom to Tabular Filters. The Bloom filter (Bloom, 1970) is the oldest and best-known filter. It works by hashing each key to random bits in a bit vector, setting them from zeros to ones. During a query, it returns a positive if all bits corresponding to a given key are set to ones and a negative otherwise. Bloom filters do not support deletes and cannot be efficiently expanded. Hence, they are typically used to represent static key sets.
Over the past 10-15 years, a new family of tabular filters emerged as an alternative to Bloom filters. A tabular filter stores a fingerprint (i.e., a hash digest) for each key in a compact hash table. It processes a query by checking if the fingerprints in a given hash slot match the fingerprint of the target key. To date, numerous tabular filters have been proposed with various nuanced performance properties on account of their different hash collision resolution strategies (i.e., cuckoo (Fan et al., 2014; Breslow and Jayasena, 2018; Wang et al., 2019), robin hood (Bender et al., 2012; Pandey et al., 2017), or two-choice (Pandey et al., 2021; McCoy et al., 2023b) hashing, etc.). Unlike Bloom filters, many Tabular filters support deletes and can associate a payload with each fingerprint. These capabilities allow tabular filters to support richer, dynamic applications (e.g., key-value stores (Dayan and Twitto, 2021; Ren et al., 2017; Conway et al., 2023; Shetty et al., 2013), large-scale k-mer analysis (McCoy et al., 2023a), duplicate detection in streams (Dutta et al., 2013), etc.).
The Need for Expandable Filters. In many applications that use filters, the data size grows over time. This establishes a need for filters that can expand along with the data. For example, many modern log-structured key-value stores employ an in-memory tabular filter to map the location of each data entry in storage (Chandramouli et al., 2018; Andersen et al., 2009; Debnath et al., 2010, 2011; Dayan and Twitto, 2021; Dayan et al., 2021; Ren et al., 2017; Conway et al., 2023). As the data size grows, the filter must expand to map a greater number of keys and their storage locations. In the networks community, expandable filters have recently been pointed out as pivotal for supporting black lists and multicast routing. In the architecture community, the 2020 ASPLOS best paper proposes to implement virtual page tables using an expandable tabular filter (Skarlatos et al., 2020). While expandable filters empower a widening range of modern applications, existing approaches exhibit performance problems.
Common Expansion Approaches. The simplest approach for expanding a filter is to reread the original data and construct a larger filter from scratch. If the data resides in storage, however, scanning it can be performance prohibitive.
Hence, the standard approach for expanding a tabular filter is to allocate a hash table with twice the capacity and to migrate every fingerprint into this larger hash table. The issue is that this requires transferring one bit from each entry’s fingerprint to become a part of its slot address to evenly distribute the fingerprints across the expanded hash table (Bender et al., 2012; Zhang et al., 2021). As a result, the fingerprints shrink as the data grows, causing the false positive rate (FPR) to increase. Eventually, the fingerprints run out of bits, at which point the filter becomes useless by returning a positive for any query.
The State of the Art: InfiniFilter. The recent InfiniFilter (Dayan et al., 2023) is a tabular filter that can grow indefinitely while better scaling the FPR. It also expands by migrating all fingerprints into a larger hash table with twice the capacity while sacrificing one bit from each fingerprint in the process. The core innovation is a new hash slot format that allows setting longer fingerprints to newer entries. On average, this stabilizes the FPR. Nevertheless, InfiniFilter suffers from two scalability concerns.
Problem 1: Performance Scalability. After the first few expansions, older fingerprints within InfiniFilter run out of bits. Such entries are referred to as void entries. A void entry cannot be uniquely mapped to a slot in an expanded filter as there is no additional fingerprint bit to transfer to its slot address. InfiniFilter addresses this problem by transferring and storing void entries across a series of smaller hash tables. The problem is that query and delete operations must probe potentially each of these hash tables, resulting in worst-case asymptotically increasing CPU overheads. Can we better scale the costs of queries and deletes?
Problem 2: FPR vs. Memory Scalability. As InfiniFilter expands, the shorter fingerprints of older entries cause the FPR to increase (i.e., at a logarithmic rate). To counteract this phenomenon, it is possible to assign increasingly long fingerprints to newer entries across expansions to cause the FPR to converge to a constant. However, this inflates the memory footprint. Is it possible to resolve or at least alleviate this contention such that the FPR vs. memory trade-off for expandable filters stays closer to that of static filters?
Aleph Filter. We propose Aleph Filter [^1] The name Aleph is borrowed from set theory, where the Aleph numbers denote different orders of infinity. , an infinitely expandable filter with constant time performance for insertions, queries, and deletes, and superior memory vs. FPR trade-offs. It achieves this by building on top of InfiniFilter in three ways.
Contribution 1: Faster Queries by Duplicating Void Entries. Aleph Filter duplicates every void entry in the expanded hash table across both slots that it could have mapped to if there were an additional fingerprint bit to sacrifice. This keeps all information about all entries, including void entries, within one main hash table. As a result, Aleph filter accesses only one hash table for any query at a cost of $O(1)$ time. We show how to tune the filter such that duplicated void entries occupy negligible space within the filter so as not to adversely impact its performance.
Contribution 2: Faster Deletes using Tombstones. Duplicating void entries within the main hash table makes it challenging to efficiently execute deletes. The reason is that potentially many duplicates must be identified and removed whenever an older entry is deleted. Aleph filter addresses this challenge by first transforming the target void entry into a tombstone in $O(1)$ time. Before the next expansion, it retrieves the original hash of the deleted void entry and uses it to identify the full collection of duplicates that need to be removed. The overhead of removing these duplicates is negligible and gets amortized as a part of the next expansion operation.
Contribution 3: Inverting the FPR vs Memory Trade-Off. In many applications, the maximum data size can be predicted in advance. We show that given a reasonable estimation of how much the data will grow, it is possible to pre-allocate slightly longer fingerprints from the onset and assign decreasing fingerprint lengths to newly inserted entries. By the time we reach the estimated data size, the filter guarantees an FPR vs. memory footprint trade-off that is on par with that of static filters. Even as the data size surpasses the estimate, Aleph Filter continues to dominate the State of the Art with respect to memory vs. FPR trade-offs that it provides.
term
definition
$N$
current filter capacity divided by initial capacity
$X$
number of expansions so far (i.e., $X=\lceil\log_{2}(N)\rceil$)
$M$
total memory used for the filter (bits / entry)
$F$
initial fingerprint length when the filter is first allocated
$h(...)$
mother hash generating function
$\alpha$
fraction of occupied slots ($10\leq\alpha<$ 1)
Table 1. Terms used throughout the paper to describe Aleph Filter and other baselines.
2. Background
This section describes Quotient Filter and InfiniFilter, on top of which we build Aleph Filter.
2.1. Quotient Filter
A quotient filter (Clerry, 1984; Dillinger and Manolios, 2009; Bender et al., 2012) is a hash table that stores a fingerprint for each inserted key. It does this by first generating a mother hash for a key consisting of $F+A$ bits using some hash function $h(...)$. The least significant $A$ bits of this mother hash represent the key’s canonical slot. The subsequent $F$ bits are the key’s fingerprint. Table 1 lists terms used throughout the paper.
A quotient filter resolves hash collisions using Robin Hood hashing (Celis et al., 1985), which is a variant of linear probing (Peterson, 1957). This means that all fingerprints mapped to a given canonical slot are stored contiguously, and they push to the right any other existing fingerprints that they collide with. A run is defined as a group of contiguous fingerprints belonging to the same canonical slot. A cluster is defined as several adjacent runs where all but the first have all been pushed to the right due to collisions.
Figure 1 Part A illustrates a quotient filter with eight slots after four insertions of Keys $V$, $Y$, $Z$ and $W$. The mother hash for each key is shown at the top. The rightmost (least significant) bits for each mother hash represent the key’s canonical address. The subsequent bits, shown in red, are the fingerprint. Keys $V$ and $Y$ share Canonical Slot 100 while Keys $Z$ and $W$ share the adjacent Canonical Slot 101. The result is a cluster consisting of two runs, each with two slots.
Metadata Bits. To keep track of the start and end of runs and clusters, a quotient filter employs three metadata bits per slot. The is__\__occupied bit indicates whether the slot is a canonical slot for at least one existing key. The is__\__shifted bit indicates whether the slot contains a fingerprint that has been shifted to the right relative to its canonical slot. The is__\__continuation bit indicates whether the slot contains the start of a new run within a cluster.
Figure 1. A Quotient filter stores a fingerprint for each key in a hash table, and it resolves has collisions by organizing fingerprints into runs and clusters. Fingerprints are illustrated in red.
In Figure 1 Part A, the is__\__occupied flag is set for Slots 100 and 101 as each of them is a canonical slot for at least one key. The is__\__shifted flag is set for Slots 101, 110, and 111 as they each contain a fingerprint that has been shifted to the right from its canonical slot. The is__\__continuation flag is set for Slots 101 and 111 as they each contain a fingerprint belonging to a run that starts to its left.
Query. A query begins at the canonical slot for the target key. If the is__\__occupied flag for this slot is set to 0, the key could not have been inserted so the query returns a negative. Otherwise, the query must find the start of the target run that might contains the sought-after key, yet this run could have been pushed to the right from its canonical slot due to collisions. To find the target run, the query first moves leftwards until reaching the start of the cluster. Along the way, it counts the number of is__\__occupied flags $c$ set to ones, each of which indicates the existence of a run in the cluster that precedes Run $r$. After reaching the start of the cluster, the query moves rightwards while skipping $c$ runs, at which point it reaches the target run. It then scans this run, returning a positive if it finds a matching fingerprint.
Figure 2. The Fingerprint Sacrifice method transfers one bit from the fingerprint to the slot address while expanding to evenly distribute the entries across the larger hash table.
Insertion. An insertion commences as a query by first finding the start of the target run for the key we wish to insert. It then adds a fingerprint to this run, pushing subsequent fingerprints in the cluster one slot to the right to clear space. This may cause more runs to join the cluster by pushing them from their canonical slots.
Delete. A delete operation also commences like a query by first finding the target run of the key we wish to remove. It then scans the run and removes the first matching fingerprint. Then, it pulls all remaining fingerprints in the cluster one slot to the left to keep the cluster contiguous. For example, Figure 1 Part B illustrates how deleting Key $Y$ pulls the second run in the cluster back to its canonical slot, causing the cluster to fragment. As with any tabular filter, it is only permissible to delete keys that we know for sure had been inserted to prevent false negatives.
False Positive Rate. A quotient filter’s FPR is $\approx\alpha\cdot 2^{-F}$, where $F$ is the fingerprint size and $\alpha$ is the fraction of the slots that are occupied. The intuition is that the target run contains on average $\alpha$ fingerprints, while the probability that each of them matches that of the sought-after key is $2^{-F}$.
Reaching Capacity. As a quotient filter exceeds $90\%$ utilization (i.e., $\alpha=0.9$), the clusters’ lengths begin to grow rapidly. This causes performance to plummet as all operations on the filter must traverse a greater number of slots. To keep performance stable and to accommodate more insertions, it is desirable to expand the filter.
Fingerprint Sacrifice. The standard approach for expanding a quotient filter is to derive the mother for every key by concatenating its canonical slot address to its fingerprint. We then reinsert the mother hash into a hash table with twice the capacity of the original one (Bender et al., 2012). Figure 2 illustrates an example. As shown, this approach transforms the least significant bit of each fingerprint to the most significant bit of its canonical slot address in order to to evenly distribute the fingerprints across the larger hash table. Hence, all fingerprints shrink by one bit in each expansion, causing the FPR to double, i.e., the FPR becomes $O(N\cdot 2^{-F})$ (Dayan et al., 2023). This method supports at most $F$ expansions, at which point all fingerprints run out of bits. We summarize the properties of this method in Row 1 of Table 2.
Figure 3. By supporting variable-length fingerprints, InfiniFilter can set long fingerprints to newer entries inserted after expanding.
2.2. InfiniFilter
InfiniFilter is a quotient filter that sets longer fingerprints to newer entries to allow expanding indefinitely while stabilizing the FPR.
Unary Padding. The central design innovation of InfiniFilter is the ability to store variable-length fingerprints. This is achieved by padding each fingerprint with a self-delimiting unary code to occupy the rest of the space in the slot. The top part of Figure 3 illustrates an instance of InfiniFilter with four occupied slots and fingerprints of lengths 3, 1, 2 and 3 bits from left to right. Fingerprints and unary codes are illustrated in red and blue, respectively.
Expansion Algorithm. As with the Fingerprint Sacrifice method, InfiniFilter transfers the least significant bit from each fingerprint to become the most significant bit of the slot address during expansion to evenly distribute existing entries across the expanded hash table. After the expansion, InfiniFilter’s slot format allows inserting longer fingerprints than the older ones that had shrunk. This is shown by the insertion of entry $Y$ in Figure 3. This entry is assigned a three bit fingerprint despite the fact that fingerprints that existed before the expansion have now all shrunk to two or fewer bits. This keeps the average fingerprint length longer.
Figure 4. InfiniFilter in the Widening Regime gradually increases the slot width across expansions to keep the false positive rate constant.
False Positive Rate. Since InfiniFilter doubles in capacity every time it expands, the fingerprints within the filter follow a geometric distribution with respect to their lengths: half of them are as long as possible, a quarter are shorter by one bit, an eighth are shorter by two bits, etc. Generally, the fingerprints created $i$ generations ago comprise a fraction of $\approx\alpha\cdot 2^{-i-1}$ of the occupied slots, while their fingerprints have a length of $F-i$ bits and hence a collision probability of $2^{-F+i}$. The FPR after $X$ expansions can be derived as a weighted average of these terms: $\sum_{i=0}^{X}2^{-F+i}\cdot\alpha\cdot 2^{-i-1}\lesssim(\log_{2}(N)+2)\cdot 2^{-F-1}\cdot\alpha$. After $\approx 2^{F}$ expansions, the FPR reaches one, at which point the filter returns a positive for any query thereby becoming useless.
Fixed-Length vs. Widening Regimes. To further stabilize the FPR, it is possible to widen slots across expansions. For example, Figure 4 illustrates increasing the slot width by one bit. Entries inserted after the expansion are set fingerprints of length four rather than three bits. By assigning fingerprints of length $F+\lceil 2\cdot\log_{2}(X+1)\rceil$ bits to new entries after the $X^{\text{th}}$ expansion, the FPR converges to a constant smaller than $2^{-F}$ (Dayan et al., 2023). The intuition is that the longer fingerprints of newer entries make their weighted contribution to the FPR increasingly small. We refer to this as the Widening Regime, in contrast to the Fixed-Width Regime described just before.
Deletes. InfiniFilter deletes a key by removing the longest matching fingerprint in the target run to prevent future false negatives. For example, Figure 5 illustrates a delete operation to Key $Z$. The target run for this key begins at Slot 1010 and consists of three slots. While the fingerprint at Slot 1010 does not match, both subsequent fingerprints, which have different lengths, do match. If we remove the shorter matching fingerprint (at Slot 1011) and it happens to belong to a different key $Y$ with a different mother hash from that of Key $Z$ (e.g., $h(Y)=011010$), future false negatives would occur when querying for Key $Y$. In contrast, removing the longest matching fingerprint (at Slot 1100) guarantees that the remaining shorter fingerprint in the run will still match the non-deleted key so that future queries to it do not result in false negatives.
Rejuvenation. In many applications (e.g., key-value stores), a positive query to a filter is followed by fetching the corresponding data entry from storage to return it to the user. In case that the key exists (i.e., a true positive), we can rehash it to derive a longer mother hash and thus rejuvenate (i.e., lengthen) the key’s fingerprint, in case it was created before the last expansion. Such rejuvenation operations help keep the FPR low. However, they are only effective when queries target older entries. Similarly to deletes, a rejuvenation operation must lengthen the longest matching fingerprint in a run to prevent false negatives.
Void Entries. After the first $F$ expansions, the oldest fingerprints in InfiniFilter run out of bits. Such entries are referred to as void entries. Both Figures 3 and 4 show the creation of a void entry at Slot 101 after the expansion. As shown, the unary code fully occupies a slot that contains a void entry. Any query that encounters a void entry in its target run returns a positive. Void entries cannot be uniquely remapped to a slot in a larger filter as there is no extra fingerprint bit to transfer to the entry’s canonical slot address.
Figure 5. InfiniFilter deletes the longest matching fingerprints in the target run to prevent future false negatives.
query/ delete/rejuv
insert
FPR
fingerprint
bits / key
max.
expansions
Fingerprint Sacrifice (Zhang et al., 2021; Bender et al., 2012)
$O(1)$
$O(1)$
$O(2^{-F}\cdot N)$
$F-O(\lg N)$
$F$
InfiniFilter (Fixed-Width Regime)
$O(\frac{\lg N}{F})$
$O(1)$
$O(2^{-F}\cdot\lg N)$
$F$
$2^{F}$
InfiniFilter (Widening Regime)
$O(\frac{\lg N}{F+\lg\lg N})$
$O(1)$
$O(2^{-F})$
$F+O(\lg\lg N)$
$\infty$
Table 2. A comparison of different techniques for expanding a filter. No existing method is scalable in terms of the costs of queries/deletes, the false positive rate, the memory footprint, and the maximum number of expansions, all at the same time.
Supporting Infinite Expansions. To support more than $F$ expansions, InfiniFilter transfers each void entry from the main hash table into a smaller secondary hash table, which has an identical structure to that of the main hash table but fewer slots. Figure 6 illustrates this process across four expansions. During the second expansion, for instance, the void entry from canonical Slot 110 of the main hash table is transferred to the secondary hash table. As 110 is also the mother hash of this entry, its least significant bit (i.e., 0) is used as a canonical slot address in the secondary hash table while its more significant two bits (i.e., 11) are used as a fingerprint.
The secondary hash table also expands when it reaches capacity. Eventually, the oldest entries in the Secondary hash table become void entries, as shown in Figure 6 by the entry at Slot 11 after the third expansion. As this point, the Secondary InfiniFilter is sealed and appended to a so-called chain of auxiliary hash tables, and a new empty secondary hash table is allocated. While this design supports an unlimited number of expansions, it slows down queries, deletes and rejuvenation operations, as they must now traverse potentially all hash tables along the chain.
Figure 6. InfiniFilter transfers void entries into a secondary hash table, which also expands when it reaches capacity. When void entries appear in the secondary hash table, it is appended to a chain of hash tables and a new secondary hash table is allocated.
3. Problem Analysis
Challenge 1: Scaling CPU Costs. The number of hash tables across which InfiniFilter stores its entries determines the CPU costs of queries, deletes and rejuvenation operations. In the Fixed-Width Regime, each hash table stores entries from across $F$ subsequent expansions, and the overall number of expansions is $\log_{2}(N)$. Hence, there are at most $O(\nicefrac{{\lg(N)}}{{F}})$ hash tables. In the Widening Regime, each hash table stores entries from across $F+O(\lg\lg N)$ subsequent expansions on average, and so the number of hash tables is $O(\nicefrac{{\lg(N)}}{{(F+\lg\lg N)}})$ (Dayan et al., 2023). We summarize these properties in Rows 2 and 3 of Table 2. Under both regimes, performance deteriorates as the data grows. Can we reduce the worst-case number of hash tables that queries, deletes, and rejuvenation operations must access?
Challenge 2: Alleviating the Memory vs. FPR Contention. As shown in Row 2 of Table 2, InfiniFilter exhibits a logarithmic FPR if we fix the number of bits per entry. Alternatively, as shown in Row 3, it exhibits a stable FPR if we increase the number of bits per entry at a doubly logarithmic rate. Is it possible to alleviate this scalability contention to achieve an FPR and memory footprint that are both on par with a static filter?
4. Aleph Filter
We introduce Aleph Filter, an expandable filter that builds on InfiniFilter to improves its scalability properties. Section 4.1 shows how Aleph Filter supports queries in $O(1)$ time by keeping and duplicating void entries within the main hash table. Section 4.2 shows analytically that the fraction of duplicated void entries stays small and therefore does not significantly impact the FPR or the maximum number of expansions that the filter supports. Sections 4.3 and 4.4 address the challenges of supporting deletes and rejuvenation operations in $O(1)$ time by lazily identifying and removing duplicated void entries. Section 4.5 shows how given a rough, conservative estimation of how much the data will grow, Aleph filter achieves superior FPR vs. memory trade-offs relative to InfiniFilter.
4.1. Fast Queries by Duplicating Void Entries
While expanding, Aleph Filter duplicates each void entry in the main hash table across both canonical slots that it could have mapped to if there were an additional fingerprint bit to sacrifice. In Figure 7, for example, there is a void entry at Slot 11 of the main hash table before the expansion. During the expansion, Aleph Filter duplicates it across Canonical Slots 011 and 111. In the next expansion, each of these duplicates will be duplicated again, resulting in four duplicates at Canonical Slots 0011, 0111, 1011, and $1$.
Figure 7. While expanding, Aleph Filter duplicates each void entry so that one duplicate would still be found in constant time when querying for the original key.
As a result of duplicating void entries, a query targeting any old entry returns a positive after one access to the main hash table. Figure 7 illustrates a query to Key $Y$, which corresponds to the void entry that was duplicated. The query visits canonical slot 111 (based on the first three bits of the key’s mother hash), finds a void entry and returns a positive. Had Key Y’s mother hash started with 011, the query would have found the other void duplicate at Slot 011 and also terminated after one hash table access. This is in contrast to InfiniFilter, which must search the smaller hash tables if the target key was inserted many expansions ago.
Another result of duplicating void entries is that any query to a non-existing key terminates after one access to the main hash table. The reason is that the main hash table contains either a fingerprint or a void entry in any existing key’s canonical slot. This means that if we do not find a matching entry in $O(1)$ time, the target entry is guaranteed not to exist. In contrast, a query to InfiniFilter must also search all the smaller hash tables to verify that the key does not exist before it can return a negative.
Hence, all queries to Aleph Filter are processed in worst-case $O(1)$ time. This is an improvement over InfiniFilter, where queries targeting non-existing or older keys must search an increasing number of hash tables as the data grows.
query/ delete/rejuv
insert
false positive rate
fingerprint
bits / key
maximum
expansions
Aleph Filter (Fixed-Width Regime)
$O(1)$
$O(1)$
$O(2^{-F}\cdot\lg N)$
$F$
$2^{F}$
Aleph Filter (Widening Regime)
$O(1)$
$O(1)$
$O(2^{-F})$
$F+O(\lg\lg N)$
$\infty$
Aleph Filter (Predictive Regime)
$O(1)$
$O(1)$
$O(2^{-F})$
$F+O(\lg|\lg(N/N_{est})|)$
$\infty$
Table 3. Aleph Filter achieves constant time performance for all operations. In addition, the new Predictive Regime significantly reduces the memory requirement given a rough estimate $N_{est}$ of how much the data will grow.
4.2. Analysis
Duplicating void entries in the main hash brings up two plausible concerns. First, do duplicated void entries significantly increase the FPR, seeing as a query that encounters a void entry immediately returns a positive? Second, can duplicated void entries fill up the main hash table to prevent the insertion of new entries? This section shows that the proportion of void entries stays moderate and therefore does not increase the FPR or decrease the maximum number of expansions that Aleph Filter supports relative to InfiniFilter.
Generational Distribution. Aleph Filter doubles in capacity during each expansion. This means that keys inserted zero, one or two expansions ago comprise approximately a half, a quarter, or an eighth of the data set, respectively, and so on. More generally, consider the Set $s_{j}$ of keys inserted in-between the $j^{\text{th}}$ and $(j+1)^{\text{th}}$ expansions, i.e., in Generation $j$. Suppose we are now in Generation $X$ ($X\geq j$), i.e., before Expansion $X+1$. Equation 1 approximates the size of Set $s_{j}$ as a fraction of the whole key set.
(1) $f(j)\approx 2^{-X+j-1}$
FPR in the Fixed-Width Regime. Consider some Generation $j$ of non-void entries (i.e., $X-j<F$) . These entries’ fingerprints each consist of $F-X+j$ bits, since they are initialized with $F$ bits and lose one bit in each expansion. The probability for a query to encounter such an entry in a given slot and falsely match its fingerprint is $\alpha\cdot f(j)\cdot 2^{-F+X-j}=\alpha\cdot 2^{-F-1}$.
Now consider a Generation $j$ of void entries (i.e., $X-j\geq F$). There are $2^{-F+X-j}$ duplicates per entry since the number of duplicates increases by a factor of two in each expansion. The fraction of slots occupied by void duplicates originating from Generation $j$ is $\alpha\cdot f(j)\cdot 2^{-F+X-j}=\alpha\cdot 2^{-F-1}$. Hence, the probability of encountering a void duplicate from Generation $j$ during a query and thus returning a false positive is $\alpha\cdot 2^{-F-1}$.
We observe that the weighted contribution of each generation of entries to the FPR is equal (i.e., $\alpha\cdot 2^{-F-1}$). Equation 2 expresses the overall FPR after $X$ expansions. As shown, Aleph Filter has approximately the same FPR as InfiniFilter in the fixed-width regime.
(2) $\small\begin{split}FPR&\approx\sum_{j=0}^{X+1}\alpha\cdot 2^{-F-1}\lesssim\alpha\cdot(log_{2}(N)+2)\cdot 2^{-F-1}\\ \end{split}$
FPR in the Widening Regime. In the Widening Regime, entries in Generation $j$ are assigned fingerprints of $\ell(j)=F+2\cdot\log_{2}(j+1)$ bits, and they lose one bit in each expansion.
Let us consider a generation Generation $j$ of non-void entries (i.e., $X-j<\ell(j)$). By Generation $X$, the fingerprints of Generation $j$ will have shrunk by $X-j$ bits to $\ell(j)-(X-j)$ bits. The probability for a query to encounter an entry from Generation $j$ in a given slot and falsely match its fingerprint is therefore $\alpha\cdot f(j)\cdot 2^{-\ell(j)+X-j}$.
Let us now suppose Generation $j$ consists of void entries (i.e., $X-j\geq\ell(j)$). The number of void duplicates for each entry is $2^{-\ell(j)+X-j}$. The probability of encountering a void duplicate originating from this generation is therefore the same expression: $\alpha\cdot f(j)\cdot 2^{-\ell(j)+X-j}$.
Equation 3 sums up this expression across all generations to derive the FPR. The derivation uses a well-known identity that the sum of the reciprocals of the square numbers (i.e., $\Sigma_{i=0}^{\infty}i^{-2}$) converges to $\nicefrac{{\pi^{2}}}{{6}}$. The result is that Aleph Filter’s FPR matches InfiniFilter’s FPR in the Widening regime.
(3) $\small\begin{split}FPR&\approx\sum_{j=0}^{X+1}\alpha\cdot f(j)\cdot 2^{-\ell(j)+X-j}\\[-5.69054pt] &\approx\alpha\cdot 2^{-F-1}\cdot\sum_{j=0}^{i}\cdot\frac{1}{(j+1)^{2}}\lesssim\alpha\cdot 2^{-F-1}\cdot\frac{\pi^{2}}{6}\lesssim\alpha\cdot 2^{-F}\end{split}$
Expansion Limit in Fixed-Width Regime. As we just saw, the fraction of slots in the filter that are occupied by void duplicates of entries created in generation $j$ is $\alpha\cdot f(j)\cdot 2^{-F+X-j}=\alpha\cdot 2^{-F-1}$. The number of generations of void entries is $X-F+1$. Hence, the fraction of slots occupied by void duplicates by the time we reach Generation $X$ ($X\geq F$) is $\gamma(X)=\alpha\cdot 2^{-F-1}\cdot(X-F+1)$. We may now ask at which point the number of duplicates takes up half the space in the filter, meaning that even after we expand, void duplicates fully take up the expanded capacity so there is no space for new insertions. We obtain it by setting $\alpha$ to one, equating $\gamma(X)$ to $\nicefrac{{1}}{{2}}$, and solving for $X$ to obtain $F+2^{F}-1$. Recall from Section 2 that the fixed-width regime only supports $2^{F}$ expansions anyways (as at this point the FPR reaches one and the filter becomes useless). Hence, Aleph Filter does not reduce the maximum number of supported expansions in the Fixed-Width Regime.
Expansion Limit in Widening Regime. As we saw, the fraction of slots in the filter occupied by void duplicates from Generation $j$ is $\alpha\cdot f(j)\cdot 2^{-FP(j)}$. Let $v$ denote the total number of generations of void entries. The overall fraction of slots occupied by void entries is $\sum_{j=0}^{v}\alpha\cdot 2^{-FP(j)}\cdot f(j)$. We note that this expression is subsumed by Equation 3 and is therefore lower than $\alpha\cdot 2^{-F}$. This establishes that the fraction of slots occupied by void entries converges to a small constant as the filter expands. The intuition is that newer entries are assigned longer fingerprints, and so it takes them increasingly longer to become void and start duplicating. As a result, Aleph Filter supports an infinite number of expansions.
4.3. Fast Deletes Using Tombstones
Duplicating void entries makes delete operations more intricate to handle. The reason is that every duplicate of a void entry that we wish to delete must be identified and removed. The challenge is do so correctly, in constant time, and without using a significant amount of additional metadata.
In the simple case that a delete operation is targeting a run with at least one matching non-void entry, the entry with the longest matching fingerprint will be removed as shown in Section 2. This section focuses on processing deletes when the only matching entries in the target run are void entries.
Figure 8. Aleph filter stores the mother hashes of all void entries along a secondary hash table and a series of auxiliary hash tables.
Identifying a Void Entry’s Duplicates. An entry becomes void when the length of the mother hash that it was assigned when it was inserted matches the logarithm base two of the number of slots in the filter. In every subsequent expansion, the number of duplicates for the entry and the number of slots in the filter both multiply by a factor of two. If the entry’s original mother hash consists of $b$ bits and the filter currently comprises $2^{k}$ slots, then the entry participated in $k-b$ expansions since it became a void entry. Therefore, it has $2^{k-b}$ void duplicates. For example, suppose the filter consists of $2^{6}$ slots and the mother hash we wish to remove is 0011. This mother hash consists of $b=4$ bits while the power of the size of the filter is $k=6$, meaning that its number of void duplicates is $42^{6-4}=$ 4.
We can also infer which canonical slots contain these void duplicates by using the mother hash of the original entry as the least significant $b$ bits of their addresses, and applying all possible permutations to the remaining $k-b$ bits. In the example just given, we use the mother hash 0011 as the least significant bits and permute the remaining two bits to obtain the following four slots addresses: 000011, 010011, 100011, 110011. To execute the delete correctly, we would need to remove a void duplicate from each one of these canonical slots. The question becomes how to efficiently store and retrieve the mother hash of any void entry so that we can identify and remove all of its void duplicates?
Secondary Hash Table. Similarly to the Chained InfiniFilter, Aleph filter adds the mother hash of any entry that turns void into a secondary hash table, which has an identical structure to that of the main hash table yet fewer slots. Figure 7 illustrates an example. Before the expansion, the secondary hash table stores only the mother hash of the void entry at Slot 11 of the main hash table. The secondary hash table then expands alongside the main hash table. After the expansion, another entry turns void at Slot 010 of the main hash table, and so its mother hash is stored in the Secondary hash table. Note that while a void entry may have multiple duplicates in the main hash table (e.g., at Slots 011 and 111), its mother hash is only stored once in the secondary hash table.
Auxiliary Hash Tables. Expanding the secondary hash table requires transferring one bit from each entry’s fingerprint to its canonical slot address in the expanded hash table. In Figure 7, for example, the entry at Slot 1 of the secondary hash table after the expansion corresponds to the entry with fingerprint 11 before the expansion, since the least significant bit of its fingerprint is repurposed as the most significant bit of its slot address. Eventually, void entries appear in the secondary hash table as well. At this point, we seal the secondary hash table and add it to a chain of auxiliary hash tables. A new empty secondary hash table is then allocated. Figure 8 illustrates the high-level workflow. This architecture is similar to that of the Chained InfiniFilter from Section 2. The core difference is that InfiniFilter traverses the secondary and auxiliary hash tables to process queries and deletes while Aleph Filter does not.
Tombstones. A delete operation commences by modifying a void entry in the canonical slot of the key to be deleted into a tombstone at the main hash table. This causes subsequent queries to the deleted key to likely return a negative (unless there is some other void entry in the slot, which would lead to a false positive). To encode a tombstone, we employ a special bit string of all ones. Figure 9 shows an example of a delete operation of Key $X$ landing at Canonical Slot 101, where the only matching entry is a void entry. The first step is changing the content of the slot from a void entry encoding (i.e., 1110) into a tombstone (i.e., $1$). As a result, subsequent queries to Key $X$ in the example will now return a negative.
Figure 9. Aleph filter deletes a void entry by first transforming it into a tombstone and storing the canonical address in a deletion queue. Before the next expansion, void duplicates are identified and removed.
Deferred Removal of Duplicates. The second step of a delete operation is to add the canonical slot of a void entry to be removed into a deletion queue, which is structured as an append-only array. Right before the next expansion (or contraction) operation, Aleph Filter pops one canonical slot address at a time from the deletion queue. It uses this address as a search key to traverse the secondary and auxiliary hash tables from largest to smallest. The search terminates as soon as it finds a matching entry. Since larger hash tables store longer mother hashes and we search the hash tables from largest to smallest, the first matching entry that is found corresponds to the longest matching mother hash. Based on this mother hash, we identify all duplicates and remove them from the main hash table [^2] Note that if many such deletes of void entries take place before the next expansion and cause the filter utilization to drop significantly below the expansion threshold again, we delay the expansion until utilization reaches the threshold again. .
In Figure 9, for example, we pop Slot 101 from the deletion queue and use it as a search key to probe the secondary hash table. We find a matching entry at Slot 1 with fingerprint of 0. We concatenating these to obtain a longest matching mother hash of 01. Since the main hash table currently has $2^{3}$ slots while the longest matching mother hash consists of two bits, we infer that the void entry has $22^{3-2}=$ 2 duplicates in the main hash table, and that their canonical slots are 001 and 101. We proceed to delete one void entry from each of these canonical slots. Finally, we also delete the longest matching mother hash from the secondary hash table.
Preventing False Negatives. Figure 10 shows a more complex example of a delete operation where there are two void entries in the target canonical slot. When we visit the Secondary hash table, we find a run containing two mother hashes of different lengths for these void entries: 00 and 000. The different lengths of these mother hashes indicate that one of the void entries has two duplicates in the filter while the other has one. There is now a question of whether to remove the entry with the single void duplicate or the one with the two void duplicates.
Let us first suppose the entry with the smaller mother hash (and hence more duplicates) is removed. If this entry happens to correspond to a different key Y with a different extended mother hash than that of key X (e.g., $h(Y)=...100$), we would get false negatives later when querying for Key $Y$. Particularly, a query for Key $Y$ would reach Slot 100, fail to find a matching entry, and return a false negative. To prevent false negatives, we must delete the void entry with the fewest duplicates. As before, this entry corresponds to the one with the longest matching mother hash. This ensures that the remaining void entry’s duplicates will still match whichever entry still exists. Hence, in Figure 10, one void entry is removed from Slot 001 of the main hash table, and the entry with Slot 0 is removed from the secondary hash table.
Figure 10. Aleph filter identifies which void duplicates to remove based on the longest matching mother hash of the deleted key.
Computational Analysis. The first two steps of a delete operation, turning a void entry into a tombstone and adding its address to the deletion queue, each cost $O(1)$ time to execute.
The cost of identifying and removing all duplicates for a given void entry is deferred and incurred right before the next expansion. To quantify this cost in the worst-case, suppose the user deletes all void entries at once. The dominating part of the cost is to remove each individual duplicate from the main filter, while the cost of retrieving every mother hash from the secondary and auxiliary hash tables is a lower-order term (i.e., $O(\nicefrac{{X}}{{F}})$). In Section 4.2, we saw that there are at most $O(2^{-F}\cdot X\cdot N)$ void duplicates in the main hash table in the Fixed-Width Regime. Each of them takes constant time to remove. Hence, the overall amount of work is $O(2^{-F}\cdot X\cdot N)$. As this work is performed after $N$ insertions, it remains sub-constant as long as $X<2^{F}$. Since this is the maximum number of supported expansions anyways, the cost of removing void duplicates is amortized constant for the filter’s whole lifetime.
In the Widening Regime, we saw in Section 4.2 that there are at most $O(2^{-F}\cdot N)$ void duplicates in the main hash table. Removing each of them after $N$ insertions entails $O(2^{-F})$ additional overhead per insertion. Hence, Aleph Filter supports deletes in constant time while being able to expand indefinitely in the Widening Regime.
Memory Analysis. The secondary and auxiliary hash tables store at most $N\cdot 2^{-F}$ mother hashes altogether. The size of these hash tables is therefore smaller by a factor of at least $2^{-F}$ from the main hash table and therefore does not take up much memory.
4.4. Fast Rejuvenation Operations
In Section 2, we saw that a rejuvenation operation rehashes a queried key after retrieving it from storage to lengthen the longest matching fingerprint in the target run. This helps reduce the FPR. In Aleph Filter, a rejuvenation operation can be trickier to handle if the only matching entry in the target run is a void entry. In this case, we must also eliminate its void duplicates.
Aleph Filter handles the case where the target run only contains matching void entries by immediately rejuvenating one void entry into the full fingerprint of the queried key. It also adds the corresponding canonical slot into a Rejuvenation Queue (similar to the Deletion Queue from Section 4.3). Just before the next expansion, Aleph filter pops one address at a time from the Rejuvenation Queue. For each address, it finds the longest matching mother hash in the Secondary or Auxiliary Hash Tables. It compares the length of this mother hash to the log of the number of slots in the main hash table to infer how many void duplicates correspond to this mother hash and what their locations are. It then removes each of these void duplicates. This process is identical to how deletes are processed with the only exception that a void duplicate need not be removed from the queried key’s canonical slot, as the void entry there has already been transformed into a full fingerprint upfront.
Figure 11 shows an example of rejuvenating a void entry at Canonical Slot 100. First, Aleph Filter transforms the void entry into a full fingerprint and adds the canonical slot address to the Rejuvenation Queue. Before the next expansion, Aleph filter pops address 100 from the Rejuvenation Queue and uses it as a search key to query the Secondary Hash Table. It finds the longest matching fingerprint of 0 at Slot 0 (the other mother hash in this run is 000, which doesn’t match 100 along the third bit). Hence, the longest matching mother hash is 00. This implies that the void entry had been duplicated twice at Canonical Slots 000 and 100 of the main hash table. However, since this is a rejuvenation operation, we know we have already replaced one void entry by a fingerprint at Canonical Slot 100, so the only remaining duplicate to remove is at Canonical Slot 000. Finally, we also remove the longest matching mother hash from the Secondary Hash Table.
Similarly to deletes, rejuvenation operations do a constant amount of work upfront while deferring and amortizing the removal of duplicates to the next expansion. Hence, their cost is $O(1)$.
Figure 11. Aleph filter rejuvenates a void entry by transforming it into a full fingerprint while adding the canonical slot to a rejuvenation queue. Before the next expansion, void duplicates are identified and removed.
4.5. Predictive Regime
So far, we have seen two complementary methods of preventing the FPR from increasing as the data grows. Rejuvenation operations lengthen fingerprints to stabilize the FPR, but they are only effective when the query workload is targeting older entries (i.e., with shorter fingerprints). On the other hand, the Widening Regime stabilizes the FPR by increasing the fingerprint length assigned to newer entries, yet the cost is a higher memory footprint comprising $F+O(\lg\lg N)$ bits per entry. This begs the question of whether there are ways to fix the FPR without relying on rejuvenation operations or using more memory. This section introduces the Predictive Regime to address this challenge.
The Predictive Regime takes as a parameter an estimation from the user of how much the data size will grow. We denote this estimate as $N_{est}$, and it is measured as the ratio between what we think the final data size will be to the initial filter capacity. Using this estimate, the Predictive Regime uses Equation 4 to assign fingerprints of length $\ell(j)$ to entries inserted at Generation $j$. The term $X_{est}=\log_{2}(N_{est})$ refers to the number of expansions before reaching the data size estimate.
At Generation 0, Equation 4 assigns larger fingerprints of length $F+\lceil 2\cdot\log(X_{est}-1)\rceil$ bits. As the data size grows towards the estimate ($1\leq j\leq X_{est}$), it assigns an equal or shorter fingerprint length to every subsequent generation. When we reach the data size estimate (i.e., $j=X_{est}+1$), the Equation assigns fingerprints of length $F$ bits, while all fingerprints assigned in previous generations will have shrunk to at most $F$ bits. Note that the max function keeps the inner part of the logarithm non-zero so that the equation is defined for all $j\geq 0$.
(4) $\ell(j)=F+2\cdot\lceil\log_{2}(\max(|X_{est}-1-j|,1))\rceil$
After surpassing the data size estimate (i.e., $j>X_{est}+1$), the Predictive Regime assumes the Widening Regime’s behavior. The absolute value function within the logarithm in Equation 4 sets increasing fingerprint lengths to subsequent generations to keep the FPR stable. Note that when we set $N_{est}=1$ (implying $X_{est}=0$), the Predictive Regime is identical to the Widening Regime from the get-go. Generally, the memory complexity for the Predictive Regime is $F+O(\lg|\lg(N/N_{est})|)$.
Visualization. Figure 12 illustrates how the fingerprint length assigned to newer entries first drops during the first $X_{est}$ expansions and then increases again afterwards. The Predictive Regime initially requires more memory than the Widening Regime. As the data grows, however, the Predictive Regime comes to improve upon the Widening Regime across the board, especially when the real data size is close to the estimate. This is a good trade-off as it is more important to require fewer bits per entry when the data is large rather than when it is small.
FPR Analysis. The analysis of the FPR with the predictive regime is similar to the analysis of the Widening Regime from Section 4.2. The contribution of any generation $0\leq j\leq X$ to the FPR is $\alpha\cdot f(j)\cdot 2^{\ell(j)-(X-j)}$. Equation 5 sums this up across all generations to obtain the overall FPR.
(5) $\small FPR=\sum_{j=0}^{X+1}\alpha\cdot f(j)\cdot 2^{\ell(j)-(X-j)}$
By plugging in $X_{est}$ for $X$ in Equation 5 and simplifying, we obtain the maximum FPR until the moment we reach the data size estimate. This turns out to be at most $2^{-F}$ by the same analysis we saw in Section 4.2 for the FPR in the Widening Regime. At this point, all fingerprints consist of at most $F$ bits. Hence, the memory vs. FPR trade-off is as good as we can hope for by the time we reach the data size estimate.
By plugging in infinity for $X$ in Equation 5 and simplifying, we obtain a maximum FPR of $2^{-F+1}$ as we surpass the data size estimate. The intuition is that the left-hand and right-hand curves of the Predictive Regime in Figure 12 each contribute an additive factor of $2^{-F}$. It is possible to assign fingerprints that are longer by one bit in advance to maintain a given FPR target as we surpass the data size estimate.
We summarize the Predictive Regime’s properties in Row 3 of Table 2. Given an accurate or even conservative estimation of how much the data size will grow, Aleph Filter provides superior FPR vs. memory trade-offs in the Predictive Regime relative to any other expandable filter.
Figure 12. Given a rough estimate of how much the data size will grow, Aleph filter requires far fewer bits per entry by the time we reach the estimate and pass it.
Figure 13. Aleph Filter exhibits faster queries while matching InfiniFilter in terms of the false positive rate, the memory footprint, and write latency.
5. Evaluation
We evaluate Aleph Filter against the Fingerprint Sacrifice (FS) method and InfiniFilter, the state-of-the-art methods for expanding filters, which were described and analyzed in Sections 2 and  3.
Implementation. We implemented Aleph Filter within InfiniFilter’s software library to enable a direct comparison. Aleph filter is built as a subclass of InfiniFilter to reuse its core machinery (e.g., for parsing slots, chaining hash tables, etc.). All baselines inherit from a common Quotient Filter class to perform hash collision resolution. The code reuse across the baselines means that any differences in their performance properties arise due to the core expansion algorithm as opposed to implementation idiosyncrasies. We employ version 11.0.16 of the Java compiler. We use xxhash (Collet, 2023) as the hash function for all baselines. For all experiments, we generate uniformly random 8-byte keys using java.util.Random. Each baseline is configured to expand when 80% of the hash table slots are occupied.
Hardware. Our system is equipped with dual Intel Xeon E5-2690v4 processors, each running at 2.6 GHz with 14 cores and two hyper-threads per core. The machine contains 512GB of RAM, 35MB of L3 cache, 256KB of L2 cache, and 32KB of L1 cache. Storage includes two 960GB SSDs and four 1.8TB HDDs, though these drives are not used in the experiments. The system runs on Ubuntu 18.04.5 LTS.
Aleph Filter Reduces Query Cost. Figure 13 shows how various cost metrics evolve as we insert data into the different baselines. All baselines are initialized with 12-bit slots. InfiniFilter and Aleph Filter employ the Fixed-Width Regime in this experiment. On each curve, one expansion occurs between any two adjacent points, indicating a doubling of the data size.
Part (A) measures query cost by averaging the latency of 10,000 queries to random non-existing keys right before the next expansion (i.e. when the filter is as full as possible). We observe that as the data size grows, the average latency gradually increases across all baselines as they outgrow the CPU caches. Nevertheless, InfiniFilter’s query cost deteriorates more rapidly since each query has to check a growing number of hash tables. The FS method cannot expand indefinitely as eventually, all fingerprints run out of bits. In contrast, Aleph Filter supports more expansions than the FS method while maintaining stabler query costs as each query checks at most one hash table.
Part (B) measures the false positive rate (FPR) for the same queries as in Part (A). The FS method exhibits a skyrocketing FPR, while InfiniFilter and Aleph Filter each exhibit a much stabler (logarithmic) FPR. Part (C) measures the memory footprint right before the next expansion. For the FS method, the number of bits per entry drops as all fingerprints become shorter by one bit in each expansion. For Aleph Filter and InfiniFilter, the memory footprint is stable at $\approx 15$ bits per entry. The reason is that while each slot consists of 12 bits, at most $80\%$ of the filter is utilized right before the next expansion. This leads to $12/0.8\approx 15$ bits per entry being used overall. Part (D) measures the average write latency while accounting for the amortized cost of expansion. As shown, Aleph Filter is at least as performant as the other baselines. Overall, Figure 13 demonstrates that Aleph Filter dominates the other methods by providing faster queries, a stabler FPR, and practically indefinite expansions, all at the same time.
Figure 14. The Predictive Aleph Filter dominates other baselines when we have a conservative estimation of how much the data will grow.
Better Trade-Offs While Growing Indefinitely. Figure 13 illustrates a similar experiment under different use-case assumptions: we envision a user that requires an FPR of at most $\approx 1\%$ while expecting the data size to grow to a maximum of $\approx 10^{6}$ entries. However, the data size in this scenario ends up growing beyond the user’s initial expectations. We initialize each baseline with the smallest memory footprint that ensures that when the data size reaches $\approx 10^{6}$ entries, the FPR is at most $\approx 1\%$. InfiniFilter and Aleph Filter in the Widening Regime are each assigned 13 bits per entry. We initialize the FS method with 30 bits per entry and Aleph filter with 22 bits per entry.
Similarly to the last experiment, Part (A) shows that the Aleph Filter outperforms InfiniFilter in terms of query cost as the data grows both in the widening and predictive regimes. At the same time, Part (D) shows that the different baselines have an approximately equal insertion latency.
Parts (B) and (C) show that the FS method is unable to meet the FPR requirement once the data size exceeds $\approx 10^{6}$ entries since the fingerprints continue to shorten in each expansion. In contrast, InfiniFilter and Aleph Filter in the Widening Regime maintain a constant FPR from the get-go and even after the data size surpasses $\approx 10^{6}$ entries. The trade-off is a slowly growing memory footprint. Aleph filter in the Predictive Regime also meets the FPR target while requiring far fewer bits per entry than in the Widening Regime when we meet the target data size. Even after we exceed the target data size, Aleph filter continues to require less memory than in the Widening regime as precisely predicted in our cost models in Figure 12. The only trade-off is that it requires a few more bits per entry from the get-go, though this is a good deal: it’s better to require more bits per entry when the data is small rather than when it is large. Overall, this experiment shows that Aleph Filter can accommodate superior FPR vs. memory trade-offs, even when we only have a rough estimate of how much the data size will grow.
Aleph Filter Supports Cheaper Deletes. Figure 15 Part (A) compares Aleph Filter to InfiniFilter in terms of delete latency. The figure illustrates two variants of Aleph Filter with greedy vs. lazy deletes. The former removes void duplicates immediately during a delete operation. The latter uses tombstones and removes void duplicates during the next expansion as described in Section 4.3. All baselines are in the Fixed-Width Regime with 12-bit slots for experimental control.
Each baseline initially consists of $2^{9}$ slots, and we perform insertions until it expands to $2^{25}$ slots. We then clone each baseline multiple times. For each clone, we measure the average latency for deleting $512$ random entries, all created in the same generation. As we move from left to right on the x-axis, we delete older entries (i.e., from earlier generations).
For all baselines, delete latency is $\approx 600$ ns when deleting recent entries from within the most recent eight generations. We observe that a delete operation costs more than queries or insertions in a quotient filter as all subsequent entries in the cluster must be shifted back by one slot to keep the cluster contiguous. As we delete older entries, the latency for InfiniFilter increases to $\approx 800$ since more hash table accesses are needed to find them. For Aleph Filter with greedy deletes, latency skyrockets as we target older entries since each entry has exponentially more duplicates in the main hash table, each of which must be removed. In contrast, for Aleph Filter with lazy deletes, latency decreases as we delete older entries since we simply replace one void entry with a tombstone (i.e., we do not need to shift the remaining fingerprints in the cluster backward by one slot). An experiment with rejuvenation operations yields a nearly identical figure to the one shown in Figure 15 Part (A), and so we omit it. This demonstrates that the use of tombstones makes Aleph filter practical for supporting dynamic data despite the myriad duplicate entries that potentially must be removed when deleting or rejuvenating an old entry.
Duplicates Removal Does not Impact Expansion Time. When Aleph filter deletes a void entry by replacing it with a tombstone, the removal of the entry’s potentially many duplicates is deferred to the next expansion. To ensure that this does not degrade expansion speed, Figure 15 Part (B) measures this overhead in both the Fixed-Width and Widening Regimes against the cost of migrating entries into an expanded hash table. We run the experiment by inserting entries into an initially small empty filter, causing it to expand multiple times. From the $10^{\text{th}}$ expansion onward, we delete the oldest remaining generation of entries before the next expansion (e.g., After Expansion 15, we delete all entries inserted in Generation 5, etc.). During the next expansion, all void duplicates for this generation of entries are removed. This experiment represents the worst-case consistent toll that deletes can inflict on expansion.
The figure shows that in the fixed-width regime, removing void duplicates takes approximately two orders of magnitude less time than the overhead of migrating entries into the expanded hash table. The reason is that the maximum number of void duplicates from each generation as a proportion of the filter size is small (i.e., $\approx 2^{-F}$ as shown in Section 4.2). In the widening regime, entries are slower to become void as they are created with increasing fingerprint lengths. Therefore, there are only a few generations of void entries to be removed. Once they are removed, all subsequent deletes occur to entries in the main hash table, and so the removal of void duplicates inflicts no toll on expansion. The experiment demonstrates that in both regimes, the removal of void duplicates is heavily amortized with respect to the cost of expansion. This ensures that insertions stay fast, no matter how many deletes targeting old entries exist in the workload.
Figure 15. Aleph Filter supports fast deletes and rejuvenation operations by using tombstones and deferring the removal of void duplicates to the next expansion.
6. Related Work
Compatibility with Existing Filters. The design principles of Aleph Filter are compatible with any underlying tabular filter that supports deriving the mother hash for each entry so that it can be reinserted into an expanded filter. Nevertheless, several reasons make the classical Quotient filter (Bender et al., 2012) particularly well-suited as a base method. First, a Quotient filter makes it trivial to derive the mother hash for each key by concatenating the fingerprint with the canonical slot address for each entry (unlike, say, Cuckoo filters (Fan et al., 2014) and their variants (Wang et al., 2019; Breslow and Jayasena, 2018)). Second, quotient filter supports deletes, which is a crucial feature for applications with dynamic data (unlike, say, XOR (Graf and Lemire, 2020), Ribbon (Dillinger et al., 2022) or Prefix filters (Even et al., 2022)).
Furthermore, we observe that while many filters have been designed to outperform the classic quotient filter when operating above $90\%$ capacity, such filters are better suited for static data sets. In the context of dynamic sets, it is more important to operate efficiently at low capacity. The reason is that by the time we reach high capacity, the filter expands, and so the filter spends most of its time at low capacity (e.g., Aleph Filter is tuned by default to expand at $80\%$ utilization, at which point it returns to $40\%$ after the expansion). Compared to other filters, a quotient filter performs particularly well at low capacity because each query incurs only one cache miss. Furthermore, clusters are small on average when operating at low capacity, meaning that the computational overhead for inserts, queries, and deletes stays modest. In contrast, a Cuckoo (Fan et al., 2014) or vector quotient filter (Pandey et al., 2021) would entail two cache misses per query, even when operating at low capacity. For these reasons, we employ quotient filter as the base method for Aleph filter, though implementing Aleph filter on top of the Rank-and-Select Quotient Filter (Pandey et al., 2017) or Crate Filter (Bercea and Even, 2020) could make for intriguing future work. In addition, applying the Aleph Filter on top of recent range filters (Goswami et al., 2014; Zhang et al., 2018; Luo et al., 2020; Knorr et al., 2022; Vaidya et al., 2022) and adaptive filters (Mitzenmacher et al., 2018; Deeds et al., 2020; Kraska et al., 2018) can also make for impactful future research.
Other Expansion Approaches. A common workaround to the filter expansion problem is to allocate additional empty filters into which more insertions can be made (Guo et al., 2006, 2009; Chen et al., 2017; Xie et al., 2007; Almeida et al., 2007; Williger and Maier, 2019). The issue with such methods is that in the worst-case, all filters need to be checked during a query. As a result, such techniques are uncompetitive with Aleph Filter in terms of query cost.
A few filter expansion approaches (Luo et al., 2019; Xie et al., 2022) form a hash ring of buckets to support elastic expansion. The issue is that queries, deletes, and insertions all become $O(\log_{2}N)$ as a search tree has to be searched to find a given entry’s bucket. In contrast, Aleph Filter performs all operations in constant time.
Yet another paradigm involves fetching a larger fingerprint for a key from storage, yet this involves issuing storage I/Os during expansion to fetch additional bits for each mother hash (Wu et al., 2021). In contrast, Aleph filter does not rely on storage I/Os to expand.
Theoretical Algorithms and Bounds. Pagh, Segev, and Wieder prove a lower bound that if we initialize a filter to constant capacity and expand it to contain $N$ keys, the filter must at some point use at least $\lg\lg N$ bits per key in addition to what is required for a filter with a fixed capacity of $N$ keys (Pagh et al., 2013). Aleph Filter’s Widening and Predictive Regimes both meet this lower bound. The former meets it after the $N$ data entries are inserted, while the latter meets it before any entries are inserted.
The idea of duplicating void entries within a filter hash table to support indefinite expansion with constant time queries can also be traced back to Pagh, Segev, and Wieder (Pagh et al., 2013). Nevertheless, their algorithm has never been implemented, and Aleph filter improves on it in four important ways. (1) Aleph filter requires $O(\lg\lg N)$ rather than $O(\lg\lg U)$ bits per key to support indefinite growth while maintaining a stable false positive rate. This is a significant difference if $U$, the universe size, is unknown or significantly larger than $N$, the data size. (2) The data structure in (Pagh et al., 2013) keeps track of the age entries by encoding a binary age counter alongside each slot for $\lg\lg N$ additional bits per key. In contrast, Aleph filter keeps track of the ages of entries using a chain of exponentially smaller hash tables that take up negligible memory. (3) The data structure in (Pagh et al., 2013) handles fingerprint collisions by storing the full keys of entries with colliding fingerprints in an auxiliary dictionary. In contrast, Aleph filter supports storing multiple identical fingerprints in the main hash table by employing an underlying quotient filter with Robin Hood hashing. This eliminates one dictionary access from the query path. (4) Aleph Filter introduces the Predictive Regime, which can drastically reduce the memory requirement.
The data structure by Liu, Yin, and Yu (Liu et al., 2020) performs constant time operations with high probability and a lower space overhead of $\lg\lg N+O(\lg\lg\lg N)$ bits per entry if $N$ and $U$ are polynomially related. This matches the leading term of the lower bound. Nevertheless, this method has not been shown to support deletes, and it does not support unbounded growth beyond a universe size of $U$.
7. Conclusion
We introduced Aleph Filter, the first practical, infinitely expandable filter data structure to simultaneously support optimal memory vs. FPR trade-offs and constant time operations.